I have to eliminate using multiplication. How would I solve these two problems?
2x+3y=6
x+2y=5
4x+5y=6
6x-7y=-20

Answer by apachemexican1(64)

2x+3y=6
x+2y=5
the first thing you must do is
cancel out the x's by adding the negative
1(2x+3y=6)
-2(x+2y=5)
you get
2x+3y=6
-2x-4y=-10
---------
y=-4
now do the exact same with y
-2(2x+3y=6)
3(x+2y=5)
------------
-4x+6y=-12
3x+6y=15
------------
-x=3
since you can't have an answer with a negative x
so you move the negative sign over
so the answer is x=-3
the points are (-3,-4)
NEXT...
4x+5y=6
6x-7y=-20
you do the same with this equation
-3(4x+5y=6)
2(6x-7y=-20)
-12x+15y=18
12x-14y=-40
------------
y=-22
do the same with y
7(4x+5y=6)
5(6x-7y=-20)
28x+35y=42
30x-35y=-100
-------------
58x=-58
now divide 58 to both sides
of the equation
58x/58=-58/58
x=-1
the points are
(-1,-22)
good luck!!!